95% CI different in proportions p1-p2

n1=70 n2 = 70

p1=.72 p2=.57

SE = .08

95% CI is _____ to _____

99% CI

n=200

p=.8

SE=.03

CI is ____ to ____

Thank you for your time.
Answer #1

Ans:

1)

Point estimate=0.72-0.57=0.15

pooled proportion=(0.72*70+0.57*70)/(70+70)=0.645

SE=SQRT(0.645*(1-0.645)*((1/70)+(1/70)))=0.08

Margin of error=1.96*0.08=0.1568

95% CI for difference in proportions p1-p2

=0.15+/-0.1568

=(-0.0068, 0.3068)

2)

SE=SQRT(0.8*(1-0.8)/200)=0.03

99% Confidence interval for p

=0.8+/-2.576*0.03

=0.8+/-0.0773

=(0.7227, 0.8773)